# Motivic cohomology of cyclic coverings

### v2

@@ Abstract
Cyclic coverings produce many examples of topologically contractible smooth affine complex varieties. In this paper, we study the motivic cohomology groups of cyclic coverings over algebraically closed fields of characteristic 00. In particular, we prove that in many situations Chow groups of cyclic coverings become trivial after tensoring with $\mathbb{Q}$. Furthermore, we can prove that the Chow groups of certain bicyclic coverings are trivial even without tensoring with $\mathbb{Q}$.
@@ 1 Introduction
The generalized Serre question on algebraic vector bundles asks whether algebraic vector bundles over topologically contractible smooth affine $\mathbb{C}$-varieties are trivial or not (cf. [, Question 6]). While it was known before that the generalized Serre question has a positive answer in dimensions $\leq 2$, it remains open in higher dimensions (cf. [, Section 5.2]). General classification results in dimension $3$ imply that a topologically contractible smooth affine complex threefold $X$ has only trivial vector bundles if and only if the Chow groups $CH^{2}(X)$ and $CH^{3}(X)$ are trivial (cf. [AF1], [KM]); the main result in [Sy] implies that all vector bundles over a topologically contractible smooth affine complex fourfold $X$ are trivial if $CH^{2}(X)$, $CH^{3}(X)$, $CH^{4}(X)$ and the Nisnevich cohomology group $H_{Nis}^{2}(X,\textbf{I}^{3})$ are trivial. Note that Chow groups are examples of motivic cohomology groups (cf. [MVW, Corollary 19.2]) and that there is a link between the group $H_{Nis}^{2}(X,\textbf{I}^{3})$ and motivic cohomology groups with $\mathbb{Z}/2\mathbb{Z}$-coefficients given by the long exact sequence in [Tot, Theorem 1.3]. A conjecture of Asok-Østvær suggests that topologically contractible smooth affine complex varieties should be stably $\mathbb{A}^{1}$-contractible (cf. [, Conjecture 5.3.11]). In view of the cohomological criteria mentioned above, this would automatically yield a positive answer to the generalized Serre question in dimensions $3$ and $4$. It is well-known that stably $\mathbb{A}^{1}$-contractible smooth affine complex varieties have the integral motivic cohomology of the base field; on the other hand, there exists a criterion for the stable $\mathbb{A}^{1}$-contractibility of a smooth affine complex variety in terms of motivic cohomology groups (cf. [DPØ, Theorem 3.1]). Altogether, it is of interest to study motivic invariants and particularly the motivic cohomology of topologically contractible smooth affine complex varieties.
@@
Definition. Let $k$ be a field of characteristic 00, let $X$ be an affine $k$-variety and let $f\in\mathcal{O}_{X}(X)\setminus\{0\}$ be a regular function and let $s>1$ be an integer. Then
@@
General Assumptions. We furthermore assume that
@@
We primarily study the motivic cohomology groups of the variety $Y_{s}$ with coefficients in a commutative ring $R$ such that $s\in R^{\times}$ and the integral motivic cohomology groups tensored with $R$. For this purpose, we first focus on the induced morphism $Y_{s}\setminus F\rightarrow X\setminus F_{0}$; by abuse of notation, we also denote this morphism by $\varphi_{s}$. This induced morphism is a finite étale Galois cover whose Galois group is the group $\mu_{s}(k)$ of $s$th roots of unity. We first prove that $\varphi_{s}$ induces an isomorphism ${H^{p,q}(Y_{s}\setminus F,R)}^{\mu_{s}(k)}\cong H^{p,q}(X\setminus F_{0},R)$ (cf. Corollary 3.7). Here ${H^{p,q}(Y_{s}\setminus F,R)}^{\mu_{s}(k)}$ denotes the subgroup of elements in $H^{p,q}(Y_{s}\setminus F,R)$ fixed by the Galois action. Motivated by this first result, we further investigate under which circumstances the Galois action on $H^{p,q}(Y_{s}\setminus F,R)$ is in fact trivial. While we are able to prove a general result on motivic cohomology groups (cf. Theorem 3.19) and, in particular, some results on the motivic cohomology groups of $Y_{s}$ with $\mathbb{Z}/n\mathbb{Z}$-coefficients for $n\geq 1$ (cf. Corollary 3.26), the main results proven in this paper concern the Chow groups of a cyclic covering (cf. Theorem 3.20):
Theorem. With the notation and assumptions above, further assume that
@@
In particular, as $H^{2i,i}(Y,\mathbb{Q})\cong CH^{i}(Y)\otimes_{\mathbb{Z}}\mathbb{Q}$ (cf. [MVW, Preface]) for $Y\in Sm_{k}$, the Chow groups of $Y_{s}$ are torsion if those of $X$ are torsion. The second assumption in the theorem also appears similarly in classical results on the singular homology of cyclic coverings (cf. [Z, Theorem 5.2]). While one needs a coefficient ring $R$ with $s\in R^{\times}$ in the theorem above, one can even prove results on the integral motivic cohomology groups $H^{2i,i}(Y_{s},\mathbb{Z})$ (i.e., on the Chow groups) of bicyclic coverings as considered in [Z, Example 6.2] (cf. Theorem 3.32):
Theorem. With the notation and assumptions above, further assume that
@@
If $k=\mathbb{C}$ and $X$ is topologically contractible in the theorem, then the variety $Y_{s,t}$ is actually known to be topologically contractible by [Z, Example 6.2]. Note that the theorem implies that the varieties $Y_{s,t}$ have only stably-trivial algebraic vector bundles. If $\dim(Y_{s,t})=3$, then general classification results even imply that all algebraic vector bundles over $Y_{s,t}$ are trivial (cf. [AF1], [KM]); if $\dim(Y_{s,t})=4$, then all algebraic vector bundles over $Y_{s,t}$ are trivial as soon as the Nisnevich cohomology group $H_{Nis}^{2}(Y_{s,t},\textbf{I}^{3})$ is trivial (cf. [Sy]). As a matter of fact, if $\dim(Y_{s,t})=4$, we are able to give a sufficient criterion for the triviality of $H_{Nis}^{2}(Y_{s,t},\textbf{I}^{3})$ and hence for the triviality of all vector bundles over $Y_{s,t}$ (cf. Theorem 3.34). Altogether, our results seem to indicate a positive answer to the generalized Serre question in dimensions 3 and 4.
Finally, we emphasize that our results make the motivic cohomology groups and particularly the Chow groups of many concrete examples of topologically contractible smooth affine complex varieties very computable in principle. As an illustration of this, we present several exemplary computations of interest in this paper (cf. Section 4); the examples include Koras-Russell threefolds of the third kind which are obtained as cyclic coverings of Koras-Russell threefolds of the first and second kinds. Koras-Russell threefolds (cf. [KR]) of the first and second kinds were studied by many mathematicians (e.g., in [HKØ] and [DF]) and have significantly shaped the research on topologically contractible smooth affine $\mathbb{C}$-varieties in the last decades.
@@ Acknowledgements
The author would like to thank Alexey Ananyevskiy, Aravind Asok, Frédéric Déglise, Adrien Dubouloz, Jean Fasel, Lucy Moser-Jauslin, Joaquín Moraga, Morgan Opie, Sabrina Pauli, Charlie Petitjean, Brian Shin, Charles Vial, Nanjun Yang and Paul Arne Østvær for helpful comments on this work. The author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project number 461453992.
@@ 2.1 Motivic cohomology
[…]
@@
[…]
@@
Representable presheaves with transfers are known to be étale sheaves with transfers (cf. [MVW, Lemma 6.2]):
Lemma 2.4.
@@
${T\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${Z\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X}$
@@ 2.2 $\mathbb{A}^{1}$-contractible varieties
For spaces $\mathcal{X},\mathcal{Y}$, we denote by $[\mathcal{X},\mathcal{Y}]_{\mathcal{H}(k)}$ the set of morphisms from $\mathcal{X}$ to $\mathcal{Y}$ in $\mathcal{H}(k)$; similarly, for pointed spaces $(\mathcal{X},x),(\mathcal{Y},y)$, we denote by $[(\mathcal{X},x),(\mathcal{Y},y)]_{\mathcal{H}_{\bullet}(k)}$ the set of morphisms from $(\mathcal{X},x)$ to $(\mathcal{Y},y)$ in $\mathcal{H}_{\bullet}(k)$.
@@
Example 2.6.
Any Nisnevich locally trivial $\mathbb{A}^{n}_{k}$-bundle $\pi:E\rightarrow X$ is an $\mathbb{A}^{1}$-weak equivalence (cf. [, Lemma 3.1.3]). In particular, this holds for any geometric vector bundle $\pi:E\rightarrow X$.
@@
If $k=\mathbb{C}$, then in low dimensions affine spaces are the unique $\mathbb{A}^{1}$-contractible smooth schemes:
Theorem 2.9.
If $char(k)=0$ and $k$ is algebraically closed, the affine line $\mathbb{A}^{1}_{k}$ is the unique $\mathbb{A}^{1}$-contractible smooth scheme of dimension $1$.
Proof.
See [AD, Claim 5.7]. ∎
Theorem 2.10.
If $k=\mathbb{C}$, the affine plane $\mathbb{A}^{2}_{\mathbb{C}}$ is the unique $\mathbb{A}^{1}$-contractible smooth scheme of dimension $2$.
Proof.
See [CR, Theorem 1.1]. ∎
@@ 2.3 Stably $\mathbb{A}^{1}$-contractible and topologically contractible varieties
The relationship between stable $\mathbb{A}^{1}$-contractibility and topological contractibility is more subtle. First of all, it follows from representability results that stably $\mathbb{A}^{1}$-contractible smooth affine varieties have the integral motivic cohomology of the base field. One has the criterion for stable $\mathbb{A}^{1}$-contractibility via Bloch’s higher Chow groups below (at least over algebraically closed fields of characteristic 00).
@@
for any smooth equidimensional quasi-projective scheme $X$ of dimension $d$ over any field $k$. As indicated above, higher Chow groups and therefore essentially motivic cohomology groups give a criterion for the stable $\mathbb{A}^{1}$-contractibility of smooth affine varieties over algebraically closed fields of characteristic 00:
Theorem 2.14.
Let $X$ be a smooth affine variety over an algebraically closed field $k$ of characteristic 00 and $x\in X$ be a closed point. If for every affine $Y\in Sm_{k}$ the projection maps $pr_{Y}:Z\times_{k}Y\rightarrow Y$ induce isomorphisms on higher Chow groups, then $(X,x)$ is stably $\mathbb{A}^{1}$-contractible.
Proof.
Follows from [DPØ, Theorem 3.1]. ∎
@@
This is [, Conjecture 5.3.11]. In dimensions $\leq 2$, this conjecture is known to be true (cf. [A, Theorem 1]); see the discussion before [, Conjecture 5.3.11] and also [, Remark 5.3.12] for further information on this conjecture.
@@
Example 2.16.
Let $k$ be any field of characteristic 00.
@@
with $m,\alpha_{2},\alpha_{3}\geq 2$ and $\langle\alpha_{2},\alpha_{3}\rangle=\mathbb{Z}$. These varieties are called the Koras-Russell threefolds of the first kind. If $k=\mathbb{C}$, the threefolds are topologically contractible. For any algebraically closed field $k$ of characteristic 00, it was proven by Hoyois-Krishna-Østvær that these varieties equipped with 00 as a basepoint are stably $\mathbb{A}^{1}$-contractible (cf. [HKØ, Theorem 4.2]); Dubouloz-Fasel even proved that Koras-Russell threefolds of the first kind are $\mathbb{A}^{1}$-contractible over any base field $k$ of characteristic 00 (cf. [DF, Theorem 1]). In case $k=\mathbb{C}$ Koras-Russell threefolds of the first kind are known to have logarithmic Kodaira dimension $\infty$- ∞ as they have non-trivial $\mathbb{C}_{+}$-actions (cf. [Z, Exercise 8.7] and [P, Introduction]), but they can nevertheless be distinguished from $\mathbb{A}^{3}_{\mathbb{C}}$ by the so-called Makar-Limanov invariant (cf. [ML]).
@@
with $b,l,\alpha_{2},\alpha_{3}\geq 2$ and $\langle\alpha_{2},b\alpha_{3}\rangle=\mathbb{Z}$. These varieties are called the Koras-Russell threefolds of the second kind and are also topologically contractible whenever $k=\mathbb{C}$. Again, Hoyois-Krishna-Østvær proved that these varieties are stably $\mathbb{A}^{1}$-contractible over any algebraically closed field of characteristic 00 ([HKØ, Theorem 4.2]). It is an open question whether they are also $\mathbb{A}^{1}$-contractible or not. As the Koras-Russell threefolds of the first kind, they have logarithmic Kodaira dimension $\infty$- ∞ (cf. [Z, Exercise 8.7] and [P, Introduction]) over $\mathbb{C}$, but can also be distinguished from $\mathbb{A}^{3}_{\mathbb{C}}$ by the so-called Makar-Limanov invariant.

### v3

@@ Abstract
Cyclic coverings produce many examples of topologically contractible smooth affine complex varieties. In this paper, we study the motivic cohomology groups of cyclic coverings over algebraically closed fields of characteristic 0. In particular, we prove that in many situations Chow groups of cyclic coverings become trivial after tensoring with $\mathbb{Q}$. Furthermore, we can prove that the Chow groups of certain bicyclic coverings are trivial even without tensoring with $\mathbb{Q}$.
@@ 1 Introduction
The generalized Serre question on algebraic vector bundles asks whether algebraic vector bundles over topologically contractible smooth affine $\mathbb{C}$-varieties are trivial or not (cf. [, Question 6]). While it was known before that the generalized Serre question has a positive answer in dimensions $\leq 2$, it remains open in higher dimensions (cf. [, Section 5.5.2]). General classification results in dimension $3$ imply that a topologically contractible smooth affine complex threefold $X$ has only trivial vector bundles if and only if the Chow groups $CH^{2}(X)$ and $CH^{3}(X)$ are trivial (cf. [AF1], [KM]); the main result in [Sy] implies that all vector bundles over a topologically contractible smooth affine complex fourfold $X$ are trivial if $CH^{2}(X)$, $CH^{3}(X)$, $CH^{4}(X)$ and the Nisnevich cohomology group $H_{Nis}^{2}(X,\textbf{I}^{3})$ are trivial. Note that Chow groups are examples of motivic cohomology groups (cf. [MVW, Corollary 19.2]) and that there is a link between the group $H_{Nis}^{2}(X,\textbf{I}^{3})$ and motivic cohomology groups with $\mathbb{Z}/2\mathbb{Z}$-coefficients given by the long exact sequence in [Tot, Theorem 1.3]. A conjecture of Asok-Østvær suggests that topologically contractible smooth affine complex varieties should be stably $\mathbb{A}^{1}$-contractible (cf. [, Conjecture 5.5.3.11]). In view of the cohomological criteria mentioned above, this would automatically yield a positive answer to the generalized Serre question in dimensions $3$ and $4$. It is well-known that stably $\mathbb{A}^{1}$-contractible smooth affine complex varieties have the integral motivic cohomology of the base field; on the other hand, there exists a criterion for the stable $\mathbb{A}^{1}$-contractibility of a smooth affine complex variety in terms of motivic cohomology groups: Indeed, if $X$ is a smooth affine complex variety and $x\in X$ is a closed point of $X$ such that for every smooth affine $\mathbb{C}$-scheme $Y$ the projection maps $pr_{Y}:X\times_{\mathbb{C}}Y\rightarrow Y$ induce isomorphisms on motivic cohomology groups, then $(X,x)$ is stably $\mathbb{A}^{1}$-contractible (cf. [DPØ, Theorem 3.1]). Altogether, it is of interest to study motivic invariants and particularly the motivic cohomology of topologically contractible smooth affine complex varieties.
@@
Definition 1.1.
Let $k$ be a field of characteristic 0, let $X$ be an affine $k$-variety and let $f\in\mathcal{O}_{X}(X)\setminus\{0\}$ be a regular function and let $s>1$ be an integer. Then
@@
General Assumptions 1.2.
We furthermore assume that
@@
We primarily study the motivic cohomology groups of the variety $Y_{s}$ with coefficients in a commutative ring $R$ such that $s\in R^{\times}$ and the integral motivic cohomology groups tensored with $R$. For this purpose, we first focus on the induced morphism $Y_{s}\setminus F\rightarrow X\setminus F_{0}$; by abuse of notation, we also denote this morphism by $\varphi_{s}$. This induced morphism is a finite étale Galois cover whose Galois group is the group $\mu_{s}(k)$ of $s$th roots of unity. Motivated by this, we then study the maps on motivic cohomology groups induced by finite étale Galois covers. For a finite étale Galois cover $\varphi:U\rightarrow V$ between smooth $k$-varieties with Galois group $G$ and of degree $s=|G|$, we prove that $\varphi$ induces an isomorphism ${H^{p,q}(U,R)}^{G}\cong H^{p,q}(V,R)$ whenever $s\in R^{\times}$ (cf. Corollary 3.5). Here ${H^{p,q}(U,R)}^{G}$ denotes the subgroup of elements in $H^{p,q}(U,R)$ fixed by the Galois action of $G$. This result applies in particular to the morphism $\varphi_{s}:Y_{s}\setminus F\rightarrow X\setminus F_{0}$. Motivated by this, we further investigate under which circumstances the Galois action on $H^{p,q}(Y_{s}\setminus F,R)$ is in fact trivial. While we are able to prove a general result on motivic cohomology groups (cf. Theorem 3.17) and, in particular, some results on the motivic cohomology groups of $Y_{s}$ with $\mathbb{Z}/n\mathbb{Z}$-coefficients for $n\geq 1$ (cf. Corollary 3.25), the main results proven in this paper concern the Chow groups of a cyclic covering (cf. Theorem 3.18):
Theorem 1.3.
With the notation of Definition 1.1 and under the General Assumptions 1.2, further assume that
@@
In particular, as $H^{2i,i}(Y,\mathbb{Q})\cong CH^{i}(Y)\otimes_{\mathbb{Z}}\mathbb{Q}$ (cf. [MVW, Preface]) for $Y\in Sm_{k}$, the Chow groups of $Y_{s}$ are torsion if those of $X$ are torsion. The second assumption in the theorem also appears similarly in classical results on the singular homology of cyclic coverings (cf. [Z, Theorem 5.2]): As $X$ is affine, a $\mathbb{G}_{m,k}$-action on $X$ is uniquely determined by its associated homomorphism of $k$-algebras $\mathcal{O}_{X}(X)\rightarrow\mathcal{O}_{X}(X)[\lambda,\lambda^{-1}]$, where $\mathcal{O}_{X}(X)[\lambda,\lambda^{-1}]$ denotes the ring of Laurent polynomials over $\mathcal{O}_{X}(X)$; then a regular function $f\in\mathcal{O}_{X}(X)$ is a quasi-invariant of weight $d\in\mathbb{Z}$ with respect to the $\mathbb{G}_{m,k}$-action if and only if this homomorphism of $k$-algebras sends $f$ to $f\lambda^{d}$. For example, when $X=\mathbb{A}^{n}_{k}=Spec(k[x_{1},...,x_{n}])$ and the $\mathbb{G}_{m,k}$-action is given by $(\lambda,x_{1},...,x_{n})\mapsto(x_{1}\lambda^{d_{1}},...,x_{n}\lambda^{d_{n}})$ for some $d_{j}\in\mathbb{Z}$, then any coordinate function $x_{j}$ is a quasi-invariant of weight $d_{j}$. While one needs a coefficient ring $R$ with $s\in R^{\times}$ in the theorem above, one can even prove results on the integral motivic cohomology groups $H^{2i,i}(Y_{s},\mathbb{Z})$ (i.e., on the Chow groups) of bicyclic coverings as considered in [Z, Example 6.2] (cf. Theorem 3.30):
Theorem 1.4.
With the notation of Definition 1.1 and under the General Assumptions 1.2, further assume that
@@
If $k=\mathbb{C}$ and $X$ is topologically contractible in the theorem, then the variety $Y_{s,t}$ is actually known to be topologically contractible by [Z, Example 6.2]. Note that the theorem implies that the varieties $Y_{s,t}$ have only stably trivial algebraic vector bundles. If $\dim(Y_{s,t})=3$, then general classification results immediately imply that all algebraic vector bundles over $Y_{s,t}$ are trivial (cf. [AF1], [KM]); if $\dim(Y_{s,t})=4$, then all algebraic vector bundles over $Y_{s,t}$ are trivial as soon as the Nisnevich cohomology group $H_{Nis}^{2}(Y_{s,t},\textbf{I}^{3})$ is trivial (cf. [Sy]). We are able to prove that $H_{Nis}^{2}(Y_{s,t},\textbf{I}^{3})$ vanishes under the assumptions of Theorem 1.4 if $\dim(Y_{s,t})=4$ and hence all vector bundles over $Y_{s,t}$ are indeed trivial (cf. Theorem 3.33). Altogether, our results seem to indicate a positive answer to the generalized Serre question in dimensions 3 and 4.
Finally, we emphasize that our results make the motivic cohomology groups and particularly the Chow groups of many concrete examples of topologically contractible smooth affine complex varieties very computable. As an illustration of this, we present several exemplary computations of interest in this paper (cf. Section 4); the examples include Koras-Russell threefolds of the third kind which are obtained as cyclic coverings of Koras-Russell threefolds of the first and second kinds. Koras-Russell threefolds (cf. [KR]) of the first and second kinds were studied by many mathematicians (e.g., in [HKØ] and [DF]) and have significantly shaped the research on topologically contractible smooth affine $\mathbb{C}$-varieties in the last decades.
@@ Acknowledgements
The author would like to thank the anonymous referee for suggesting changes which greatly improved the exposition of the paper. The author would like to thank Alexey Ananyevskiy, Aravind Asok, Frédéric Déglise, Adrien Dubouloz, Jean Fasel, Lucy Moser-Jauslin, Joaquín Moraga, Morgan Opie, Sabrina Pauli, Charlie Petitjean, Brian Shin, Charles Vial, Nanjun Yang and Paul Arne Østvær for helpful comments on this work. The author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project number 461453992.
@@ 2.1 Motivic cohomology
A presheaf with transfers is an additive functor $Cor_{k}^{op}\rightarrow Ab$. Analogously, for any commutative ring $R$, a presheaf with transfers of $R$-modules is an additive functor $Cor_{k}^{op}\rightarrow\textit{R-mod}$.
@@
Note that any presheaf with transfers (of $R$-modules) can be restricted to a presheaf $Sm_{k}^{op}\rightarrow Ab$ (to a presheaf $Sm_{k}^{op}\rightarrow\textit{R-mod}$). In particular, one has a natural definition for sheaves with transfers:
@@
Representable presheaves with transfers are known to be étale sheaves with transfers:
Lemma 2.4 ([MVW, Lemma 6.2]).
@@
[…]
@@ 2.2 $\mathbb{A}^{1}$-contractible varieties
For spaces $\mathcal{X},\mathcal{Y}$, we denote by $\mathcal{X},\mathcal{Y}]_{\mathcal{H}(k)}$ the set of morphisms from $\mathcal{X}$ to $\mathcal{Y}$ in $\mathcal{H}(k)$; similarly, for pointed spaces $(\mathcal{X},x),(\mathcal{Y},y)$, we denote by $(\mathcal{X},x),(\mathcal{Y},y)]_{\mathcal{H}_{\bullet}(k)}$ the set of morphisms from $(\mathcal{X},x)$ to $(\mathcal{Y},y)$ in $\mathcal{H}_{\bullet}(k)$.
@@
Example 2.6 ([, Lemma 5.3.1.3]).
Any Nisnevich locally trivial $\mathbb{A}^{n}_{k}$-bundle $\pi:E\rightarrow X$ is an $\mathbb{A}^{1}$-weak equivalence. In particular, this holds for any geometric vector bundle $\pi:E\rightarrow X$.
@@
If $k$ is a field of characteristic 0, then in low dimensions affine spaces are the unique $\mathbb{A}^{1}$-contractible smooth schemes:
Theorem 2.9 ([AD, Claim 5.7]).
If $char(k)=0$, the affine line $\mathbb{A}^{1}_{k}$ is the unique $\mathbb{A}^{1}$-contractible smooth scheme of dimension $1$.
Theorem 2.10 ([CR, Theorem 1.1]).
If $char(k)=0$, the affine plane $\mathbb{A}^{2}_{k}$ is the unique $\mathbb{A}^{1}$-contractible smooth scheme of dimension $2$.
@@ 2.3 Stably $\mathbb{A}^{1}$-contractible and topologically contractible varieties
The relationship between stable $\mathbb{A}^{1}$-contractibility and topological contractibility is more subtle. First of all, it follows from representability results that stably $\mathbb{A}^{1}$-contractible smooth affine varieties have the integral motivic cohomology of the base field. One has the criterion for stable $\mathbb{A}^{1}$-contractibility via Bloch’s higher Chow groups below (at least over algebraically closed fields of characteristic 0).
@@
for any smooth equidimensional quasi-projective scheme $X$ of dimension $d$ over any field $k$. As indicated above, higher Chow groups and therefore essentially motivic cohomology groups give a criterion for the stable $\mathbb{A}^{1}$-contractibility of smooth affine varieties over algebraically closed fields of characteristic 0:
Theorem 2.14 ([DPØ, Theorem 3.1]).
Let $X$ be a smooth affine variety over an algebraically closed field $k$ of characteristic 0 and $x\in X$ be a closed point. If for every affine $Y\in Sm_{k}$ the projection maps $pr_{Y}:X\times_{k}Y\rightarrow Y$ induce isomorphisms on higher Chow groups, then $(X,x)$ is stably $\mathbb{A}^{1}$-contractible.
@@
In dimensions $\leq 2$, this conjecture is known to be true (cf. [A, Theorem 1]); see the discussion before [, Conjecture 5.5.3.11] and also [, Remark 5.5.3.12] for further information on this conjecture.
@@
Example 2.16 ([HKØ, (3.1), (3.2)]).
Let $k$ be any field of characteristic 0.
@@
with $m,\alpha_{2},\alpha_{3}\geq 2$ and $\langle\alpha_{2},\alpha_{3}\rangle=\mathbb{Z}$. These varieties are called the Koras-Russell threefolds of the first kind. If $k=\mathbb{C}$, the threefolds are topologically contractible. For any algebraically closed field $k$ of characteristic 0, it was proven by Hoyois-Krishna-Østvær that these varieties equipped with 0 as a basepoint are stably $\mathbb{A}^{1}$-contractible (cf. [HKØ, Theorem 4.2]); Dubouloz-Fasel even proved that Koras-Russell threefolds of the first kind are $\mathbb{A}^{1}$-contractible over any base field $k$ of characteristic 0 (cf. [DF, Theorem 1]). In case $k=\mathbb{C}$ Koras-Russell threefolds of the first kind are known to have logarithmic Kodaira dimension -$\infty$ as they have non-trivial $\mathbb{C}_{+}$-actions (cf. [Z, Exercise 8.7] and [P, Introduction]), but they can nevertheless be distinguished from $\mathbb{A}^{3}_{\mathbb{C}}$ by the so-called Makar-Limanov invariant (cf. [ML]).
@@
with $b,l,\alpha_{2},\alpha_{3}\geq 2$ and $\langle\alpha_{2},b\alpha_{3}\rangle=\mathbb{Z}$. These varieties are called the Koras-Russell threefolds of the second kind and are also topologically contractible whenever $k=\mathbb{C}$. Again, Hoyois-Krishna-Østvær proved that these varieties are stably $\mathbb{A}^{1}$-contractible over any algebraically closed field of characteristic 0 ([HKØ, Theorem 4.2]). It is an open question whether they are also $\mathbb{A}^{1}$-contractible or not. As the Koras-Russell threefolds of the first kind, they have logarithmic Kodaira dimension -$\infty$ (cf. [Z, Exercise 8.7] and [P, Introduction]) over $\mathbb{C}$, but can also be distinguished from $\mathbb{A}^{3}_{\mathbb{C}}$ by the so-called Makar-Limanov invariant.